OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2?

OpenStudy (anonymous):

@robtobey

OpenStudy (jdoe0001):

notice the sequence 5, 2 how much was added/subtracted to "5" to get 2?

OpenStudy (anonymous):

3?

OpenStudy (jdoe0001):

hmm +3?

OpenStudy (anonymous):

no -3

OpenStudy (jdoe0001):

then THAT is the "common difference" or "d" in the equation for an arithmetic sequence, thus \(\bf a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1){\color{blue}{ d}}\implies a_{\color{brown}{ n}}=5+({\color{brown}{ n}}-1){\color{blue}{ (-3)}}\)

OpenStudy (anonymous):

an = 5 − 3(n − 1); all integers where n ≥ 1 or an = 5 − 3(n − 1); all integers where n ≥ 0

OpenStudy (jdoe0001):

well... the "n" a term.... and a term usually starts off at 1 onwards..... so \(\bf n \ge 1\)

OpenStudy (anonymous):

thank you! can you help with some more? ill post them

OpenStudy (jdoe0001):

yw sure, post away